The table below shows the weekly cost of producing cowboy hats.

Complete the table by filling in the missing values.

Instructions: Round your answers to one decimal place.

Cowboy Hat Production Costs

Output Total Fixed Cost (dollars) Total Variable Cost (dollars) Total Cost (dollars) Average Fixed Cost (dollars) Average Variable Cost (dollars) Average Total Cost (dollars)
0 $1,500 $0 $1,500
10 200 $   $   $170
20 1,860 18 93
30 560 50 18.7
40 2,300 57.5
50 1,100 22
Answer #1
Output TFC TVC=TC-TFC TC= TFC+TVC=ATC(Q) AFC=FC/Q AVC=TVC/Q ATC=TC/Q
0 1500 0 1500 - - -
10 1500 200 1700 1500/10=150 200/10=20 1700/10=170
20 1500 (1860-1500)=360 1860 1500/20=75 18 93
30 1500 560 (1500+560)=2060 1500/30=50 18.7 2060/30=68.7
40 1500 (2300-1500)=800 2300 1500/40=37.5 800/40=20 57.5
50 1500 1100 (1500+1100)=2600 1500/50=30 22 2600/50=52